Part A What is the reactance of a 2.40 μF capacitor at a frequency of 80.0 Hz ?

Part B

What is the reactance of a 2.40 μF capacitor at a frequency of 1.80 MHz ?
Answer #1

Xc1 = 1/(ω1*C) = 10^6/(2*3.1416*80*2.40) = 828.93 ohm
Xc2 = Xc1*f1/f2 = 828.93*80/(1.8*10^6) = 0.036841 ohm